# From Internal Conflict to Contextual Adaptation of Language Models

### v1

@@ 3.3 Coherent Persuasion Score
The degree of context-memory conflict can be quantified by measuring a context’s efficacy in shifting the LM’s answer. In previous work, persuasion scores were proposed to see to what degree context changes a LM’s answer Du et al. (2024). However, the existing persuasion score focuses on the first token of the single generated answer which has been regarded as insufficient to represent the whole generated sequence Wang et al. (2024). There also exists the problem of uncertainty during the generation to solely rely on a single forward pass, which may also stem from syntactical uncertainties, which we are irrelevant for our use-case.
@@
To overcome the LM’s brittleness, we propose the novel Coherent Persuasion (CP) score. We incorporate a multiple sample approach and semantically group the samples to create distributions for comparison. Specifically, we adapt the generation and grouping process as in Kuhn et al. (2023) into the CP score.
@@
To calculate the CP score of the $c_{i}$, we gather two lists of answers, $A_{q_{i}}$ and $A_{c_{i}\&q_{i}}$, with two different inputs, which correspond to the question only input $[P;q_{i}]$, and the question with context input $[P;c_{i};q_{i}]$. Then, we create the groups $G_{q_{i}}=[g_{1},g_{2},...,g_{r},...,g_{R}]$ and $G_{c_{i}\&q_{i}}=[g_{1},g_{2},...,g_{u},...,g_{U}]$ from $A_{q_{i}}$ which has $R$ number of generated answers and $A_{c_{i}\&q_{i}}$ which has $U$ number of generated answers, respectively. The CP score is obtained by averaging the distance from the probability distribution $p$ of $G_{c_{i}\&q_{i}}$ to $p$ of $G_{q_{i}}$. In detail, for the probability distribution of $g_{r}$, denoted as $p_{g_{r}}$, is obtained as:
${}p_{g_{r}}=\frac{1}{W}\sum_{w=1}^{W}p_{a_{w}}$ (3)
@@
where the $g_{v}$ has $W$ generated answers and $p_{a_{w}}$ is the averaged softmax probability distribution of all the tokens in the answer $a_{w}$. The $p_{g_{u}}$ is also calculated following Equation 3. Our final proposed CP score can be acquired by Equation 4.
${}CP(c_{i})=\frac{1}{|R|\times|U|}\sum_{r=1}^{R}\sum_{u=1}^{U}kl\_div(p_{g_{r}},p_{g_{u}})$ (4)
@@
As the CP score measures the distance between the probability distributions of $A_{c_{i}\&q_{i}}$ and $A_{q_{i}}$, it tells us how much the LM’s probability distribution is swayed by the context, while ensuring that small changes in syntactic representation or mismatches from first-token probabilities are not included.
4 DynamicQA
To evaluate our methods proposed in §3, we present a dataset of 11,378 question-answer pairs that features the dynamicity of facts (see Figure 1). The dynamicity of a fact is difficult to annotate, so we use naturally occurring proxy scores as a reflection of the aspect. The question and answer pairs, alongside their temporality and disputability scores, are sourced from Wikidata and Wikipedia, which has not previously been provided in related datasets. We approximate temporality via the number of edits (§4.1) and disputability via the number of reversions (§4.2). For each question and answer pair, we obtain answer-specific context snippets from Wikipedia, alongside their popularity scores. For static and temporal facts, we also create ‘counter-memory’ from these snippets using similar object replacements, to simulate realistic knowledge conflict scenarios.
4.1 Temporality
We define the temporality of a triplet via the number of edits on Wikidata for the given subject and relation as a proxy. We initialize our dataset from PopularQA Mallen et al. (2023), which is a collection of 14k question-answer pairs sampled from Wikipedia, with a popularity score determined by the monthly Wikipedia page views of the subject and object of the triplets. This popularity score is often used as a proxy for the triplet’s prevalence in the unsearchable pre-training corpora of models Mallen et al. (2023); Fierro et al. (2024). Given the sampling method, this dataset has a long-tailed distribution, meaning most triplets have very low popularity and low temporality. We identify relevant snippets from the triplet’s subject current Wikipedia page that mention the triplet’s object as context to provide to the LM. To simplify replacement of the object for counter-memory context at test-time, we replace the object in context with a replacement token. We discard 2k pairs where we cannot find the intended object and relation mentioned on the subject’s Wikipedia page. We identify relevant replacement entities using the most similar entity for the original object as identified using the Wembedder tool Årup Nielsen (2017). All facts with more than 1 edit are labeled as ‘temporal’ facts, though they vary in their degree of temporality (with a maximum score of 23 edits). This leaves us with 2495 questions. We randomly subsample 2500 of the remaining ‘static’ questions (triplets with 0 edits on Wikidata).
@@ 4.1 Temporality
Table 1: Number of instances for each type of facts
4.2 Disputability
Disputable facts are facts that could be changed depending on the viewpoint. To construct the question and answer pairs containing disputability, we utilize the collection of articles from Wikipedia that are regarded as controversial. [^1] https://en.wikipedia.org/wiki/Wikipedia:List_of_controversial_issues Controversy, by definition, is inherently disputable since it arises from the existence of multiple viewpoints, each supported by sufficient evidence. Rad and Barbosa (2012); Vuong et al. (2008). We are the first to convert the controversies that appear in the edit history into a QA dataset for an LM.
To identify the disputable facts from the articles, we look into reverted edit logs Sumi et al. (2011). Among the edit logs $\{e_{0},e_{l-1},e_{l},e_{l+1},...,e_{j}\}$ on an article, two consecutive logs, $e_{l}$ and $e_{l+1}$ are selected as a pair of reverted edit logs if the texts of $e_{l-1}$ and $e_{l+1}$ are identical. Then, by measuring the edit distance between $e_{l}$ and $e_{l+1}$, we obtain $t$ number of modification pairs $\{(m_{l,1},m_{l+1,1}),(m_{l,s},m_{l+1,s}),...,(m_{l,t},mt_{l+1,t})\}$. The pair is selected as a potential disputable pair if replacing $m_{l,s}$ to $m_{l+1,s}$ in $e_{l}$ yields the same text with the $e_{l+1}$.
Since a reverted edit does not necessarily imply disputability of the fact, we apply a few rules to filter out inappropriate modification pairs, to account for cases of vandalism, paraphrasing, or synonyms. For vandalism, we remove the pair if one of the users involved in that edit did not disclose their identity, for example, anonymous users or if the user ID is an IP address. For the modification pairs that are synonyms or paraphrasing, we feed the pairs to the semantic similarity model (§A.3) and filter the pairs whose similarity scores are bigger than 0.98. The selected pairs of reverted edit logs and modification pairs serve as contexts and answers in our dataset.
Questions are generated using a LM [^2] https://huggingface.co/meta-llama/Meta-Llama-3-8B-InstructA.2) by providing the context $e_{l}$ and the answer $m_{l,s}$ within the prompt. After obtaining the question and answer pair with the context, three annotators performed a manual annotation of the dataset to ensure the quality of the dataset; two annotators annotated each datapoint, and conflicts were resolved by the third annotator. We obtain a Krippendorf’s alpha of 0.44 and provide further details in Appendix A.3. For the real-world proxy score of the disputability, we count the number of reversions on $(e_{l},e_{l+1})$. As a result, we obtain 694 questions with two possible disputable answers and its real-world proxy score (Table 1).
Accuracy ($\uparrow$) Uncertainty ($\downarrow$) Persuasion ($\uparrow$)
@@ 4.2 Disputability
Table 2: We present the average accuracy, uncertainty and persuasion score of the three investigated models.
Figure 2: The change in semantic uncertainty given context across all tested models.
5 Experiments
5.1 Setup
We use DynamicQA to assess the performance of three recent and similarly sized state-of-the-art LMs: Mistral-7B-Instruct-v0.1 (Mistral) Jiang et al. (2023), Llama-2-7b-chat-hf (Llama-2) Touvron et al. (2023), and Qwen2-7B-Instruct (Qwen2) Qwen (2024). To minimize the effect of the confounding factors, inferences are done in zero-shot manner, and, to obtain better generation results, instruction-tuned LMs are chosen. To obtain the model’s parametric knowledge ($y_{i1}$), we first query the model for each $q_{i}$ without any additional context. We then query the model provided two forms of context: one is the unperturbed context $c_{o}$ and the other is the unseen replacement $c_{c}$. In the case of disputable instances, the choice of $c_{o}$ and $c_{c}$ is arbitrary, as both options are equally likely. For each query, we obtain the accuracy and semantic uncertainty (§3.2). We calculate the CP score for each $q_{i}$ across all queries (§3.3). We calculate the accuracy of each model on each partition as
$acc=\frac{\sum_{i=1}^{N}RougeL(a_{i},y_{i})>0.3}{N}$
where $a_{i}\in\{a_{c},a_{o}\}$, when context is provided, and is otherwise $a_{o}$. Given the inherent subjectivity of questions within the Disputable dataset, we do not calculate accuracy for that partition without context provided. For generating samples, we set $K=10$ and calculate accuracy values using the most likely response as determined via greedy search.
5.2 Analysis of Semantic Uncertainty and Coherent Persuasion Scores
We assess our semantic uncertainty score in relation to the dynamicity of facts to determine if it is a reliable indicator of intra-memory conflict. Furthermore, we look at the interaction of our intra-memory conflict indicators (dynamicity scores, and semantic uncertainty) with the CP score to elucidate the interaction between our two investigated forms of knowledge conflict.
We identify two model behaviours of interest: Persuaded instances are instances where the model is persuaded by the provided context, meaning that $y_{i1}\neq a_{i}$ but $y_{i2}=a_{i}$. Stubborn instances are instances where the model is impassive to the provided context, meaning $y_{i1}=y_{i2}$ and $y_{i1}\neq a_{i}$. We compare the obtained persuasion and uncertainty scores across the entire dataset and also highlight the behaviour of the persuaded and stubborn instances (§6.1, §6.2). Furthermore, we identify datapoints of interest. In the temporal dataset, we identify the top 100 most temporal datapoints (i.e. top-100 most edits), and the top 100 most popular datapoints (i.e. top-100 subject pageviews). We assess for differential distributions in uncertainty and persuasion for these remarkable datapoints (§6.3, §6.4) to assess for the relevance of temporality and popularity in retrieval-augmented generation and QA Mallen et al. (2023).
To further investigate the influence of the various potential predictors of an instance’s persuasion, we implement a linear regression model across all instances, where we take as independent variables the popularity of the subject and object, the number of edits to the fact, and the semantic uncertainties before and after context is provided (§6.5). We standardise each parameter with a $z$-score transformation for interpretable comparison. The provided $\hat{\beta}$ values reflect the magnitude of each predictor’s effect on the dependent variable, the persuasion score, and the $p$-value denotes the statistical significance of the effect.
@@ 6.1 General Performance
Table 2 shows the performance of our three investigated models across the three partitions of our dataset with and without additional context. Table 1 also shows the percentage of stubborn and persuaded instances. We see that accuracy is typically greatest with Llama-2 across all partitions. We also see that Llama-2 has the greatest degree of persuasion (in both persuasion score and % persuaded) across all partitions. However, Llama-2 shows the greatest uncertainty across all instances. Typically, uncertainty decreases with context; however, Mistral and Qwen2 show a slight increase in uncertainty for the Static partition of the dataset. Furthermore, we typically see greatest accuracy and uncertainty on the Static dataset, though the difference in uncertainty between the Static and Temporal partitions are not stark.
@@
These results are somewhat similar to contemporary work Fierro et al. (2024), who also find the greatest accuracy with Llama-2; however, they also find the highest confidence with this model, though they presumably use the output of softmax layer which is more vulnerable to syntactic uncertainties Kuhn et al. (2023). Furthermore, they find reduced update accuracy for Llama-2 relative to other models across mutable and immutable facts. In contrast, we surprisingly find that manipulations to static, immutable facts are most easily accepted by the model in context. In contrast, the temporal and disputable facts, which are expected to have greater variability in the model’s training dataset, and thus higher intra-memory conflict, have a greater proportion of stubborn instances across all models. We expect this stems from our differences in dataset collection, as Fierro et al. (2024) determine mutability via relationships, whereas we verify the ‘mutability’ of our facts with our proxy scores. Ultimately, our results raise concerns for the efficacy of retrieval-augmented text generation, as it appears that the most commonly adapted facts are those that are the most difficult to update with context.
6.2 Model-Level Differences
@@
In Figure 2, we show model-level differences in behaviour with provided context. We typically see a decrease in uncertainty given context. In the case of Mistral, stubborn instances have the greatest decrease in uncertainty, whereas persuaded instances have a minimal change. Llama, in contrast, has the smallest decrease in uncertainty for stubborn instances. Qwen2 has minimal change in uncertainty for both persuaded and stubborn instances. We further investigate differences between persuaded and stubborn instances in the next section. For the main paper, we present results for Llama, due to its high performance, on the temporal partition of the dataset, as this partition allows for easy comparison of the effect of dynamicity on model performance. We provide all the additional graphs for the remaining models and partitions in Appendix B.
6.3 Interaction of Persuasion and Uncertainty
Figure 3: The instance-level relationship between the persuasion and uncertainty with and without context. Persuaded and stubborn instances are denoted with orange and black markers.
@@ 6.3 Interaction of Persuasion and Uncertainty
Furthermore, we see an interesting behaviour given the introduction of context: While the majority of instances given context, for Llama, decrease in uncertainty, there is a small cluster of datapoints that do not show this behaviour. This dichotomy of behaviour is also seen in other models (See Appendix B), even though the direction of uncertainty change differs (as well as number of clusters). The cause for this behaviour is unclear, though we do investigate other potential factors (like temporality, popularity and keywords) in Appendix C, where we do not find a definitive pattern.
6.4 Interaction of Temporality and Uncertainty
In Figure 4, we show the distribution of highly popular and temporal instances between persuaded and stubborn instances. We can see that, though highly popular values typically show lower semantic uncertainty without context, they are not overly represented among the stubborn instances. Highly temporal values, on the other hand, typically show greater semantic uncertainty without context. We also do not see a particularly large decrease in the semantic uncertainty of highly temporal facts. Therefore, the semantic uncertainty measure does not seem to be an accurate indicator of intra-memory conflict, or aleatoric uncertainty, as previously used in other work Gao et al. (2024). This suggests that other factors may contribute to semantic uncertainty, outside of exclusively intra-memory conflict.
Figure 4: The distribution of semantic uncertainty in persuaded and stubborn models, with reference to all instances. The top-100 temporal and popular values are denoted.

### v2

@@ 4.3 Coherent Persuasion Score
We posit that the decreased context utilization in the case of highly popular facts Mallen et al. (2023); Du et al. (2024) owes to the greater likelihood attributed to the learned answer in the unconditioned output distribution (given $x_{i;q}$). Therefore, greater effort, either via parameter updates or contextual knowledge, is required to shift this output distribution to match the new answer provided in the context ($a_{c}$) in the conditioned instance (given $x_{i;q,c}$, Figure 1b). This is further exacerbated when there is a superposition of competing answers (Figure 1c). The amount of effort that must be enacted on the model parameters to output $a_{c}$ can be approximated by the loss, which also reflects the LM’s perplexity to the output. We also measure the observed magnitude of this shift in output distribution via our novel Coherent Persuasion ($CP$) score. This score quantifies that actual efficacy of the context in shifting an LM’s output distribution.
@@
Previously, persuasion scores were introduced to assess to what degree context changes an LM’s answer Du et al. (2024). However, the existing persuasion score focuses on the first token of the single generated answer, which has been shown to be insufficient to represent the entire generated sequence Wang et al. (2024). There also exists the problem of uncertainty during the generation of a single forward pass, which may also stem from syntactical uncertainties, which are irrelevant to our use case.
@@
To overcome the LM’s brittleness, we propose the novel Coherent Persuasion ($CP$) score. We incorporate a multiple-sample approach and semantically group the samples to create answer distributions for both $x_{i;q}$ and $x_{i;q,c}$ and compare the two. To create answer distributions, we adapt the generation and semantic grouping process as in Kuhn et al. (2023) into the $CP$ score.
@@
To calculate the $CP$ score of our context $c_{i}$, we gather two lists of answers, $Y_{i;q}$ and $Y_{i;q,c}$, with two different inputs, which correspond to the question only input $x_{i;q}$, and the question with context input $x_{i;q,c}$. Then, we create the semantically similar groups $G_{i;q}=[g_{1},g_{2},...,g_{r},...,g_{R}]$ and $G_{i;q,c}=[g_{1},g_{2},...,g_{u},...,g_{U}]$ from $Y_{i;q}$ which has $R$ groups of generated answers and $Y_{i;q,c}$ which has $U$ groups of generated answers, respectively. The $CP$ score is then obtained by averaging the divergence from the probability distribution $p$ of $G_{i;q,c}$ to $p$ of $G_{i;q}$. In detail, the output probability distribution of $g_{r}$ and $g_{u}$, is obtained as:
${}p_{g_{w}}=\frac{1}{W}\sum_{w=1}^{W}p_{y_{w}}$ (3)
@@
where $W\in\{R,U\}$ and $p_{y_{w}}$ is the averaged softmax probability distribution of all the tokens in the answer $a_{w}$. Our final proposed $CP$ score can be acquired by Equation 4.
${}CP(c_{i})=\frac{1}{|R|\times|U|}\sum_{r=1}^{R}\sum_{u=1}^{U}KL(p_{g_{r}},p_{g_{u}})$ (4)
@@
As the $CP$ score approximates the distance between the probability distributions of $Y_{i;q,c}$ and $Y_{i;q}$, it tells us how much the LM’s probability distribution is swayed by the context, while ensuring that small changes in syntactic representation or mismatches from first-token probabilities are not included. It is therefore more ‘coherent’ in that it considers (1) the entirety of the LM’s output and (2) only the semantic divergence in output.
5 Experiments
5.1 Setup
We use DynamicQA to assess the performance of three recent and similarly sized state-of-the-art LMs: Mistral-7B-Instruct-v0.1 (Mistral, Jiang et al. (2023)), Llama-2-7b-chat-hf (Llama-2, Touvron et al. (2023)), and Qwen2-7B-Instruct (Qwen2, Yang et al. (2024)). To minimize the effect of the confounding factors, inferences are done in zero-shot manner, and, to obtain better generation results, instruction-tuned LMs are chosen. To obtain the model’s parametric knowledge ($y_{i;q}$), we first query the model for each $q_{i}$ without any additional context. We then query the model provided two forms of context: one is the unperturbed context $c_{o}$ and the other is the unseen replacement $c_{c}$. In the case of disputable instances, the choice of $c_{o}$ and $c_{c}$ is arbitrary, as both options are equally likely. For each query, we obtain the accuracy and semantic entropy (§4.2). We calculate the $CP$ score for each $q_{i}$ across all queries (§4.3). We calculate the accuracy of each model on each fact type as
$acc=\frac{\sum_{i=1}^{N}RougeL(a_{i},y_{i})>0.3}{N}$
where $a_{i}\in\{a_{c},a_{o}\}$, when context is provided, and is otherwise $a_{o}$. Given the inherent subjectivity of questions within the disputable portion of the dataset, we do not calculate accuracy for that partition without the context provided. For generating samples, we set $K=10$ and calculate accuracy values using the most likely response as determined via greedy search.
@@ 5.1 Setup
Table 1: The number of collected questions and instances in DynamicQA3) for each fact type. We also report general model behaviour (i.e. percentage of persuaded instances given context), as further described in §5.2.
Accuracy ($\uparrow$) Semantic Entropy ($\downarrow$) Coherent Persuasion Score ($\uparrow$)
@@
Table 2: The average accuracy, Semantic Entropy ($SE$; §4.2) and Coherent Persuasion ($CP$; §4.3) score of our models. We bold the best values per column, with and without context. Given the inherent subjectivity of the Disputable facts, we do not show accuracy without context.
5.2 Analysis
We first look into general differences in model performance on static and dynamic facts, by comparing changes in accuracy, entropy and persuasion.
We identify the difficulty of updating dynamic facts by identifying two model behaviours of interest: Persuaded instances are instances where the model is persuaded by the provided context, meaning that $y_{i;q}\neq a_{c}$ but $y_{i;q,c}=a_{c}$. Stubborn instances are instances where the model is impassive to the provided context, meaning $y_{i;q}=y_{i;q,c}$ and $y_{i;q}\neq a_{c}$. These are exceptional instances of RAG context usage failure. We compare the obtained $CP$ and $SE$ scores across the entire dataset and also highlight the behaviour of the persuaded and stubborn instances (§6.1). To further explain the difficulty of retrieval-guided model updates, we observe the differences in loss (§6.2).
To identify factors contributing to persuasion (or RAG context usage success), we look at three characteristics of interest on the temporal dataset: semantic entropy, temporality (the number of edits), and popularity (approximated by pageviews). We analyse the Pearson correlation between these three values and our $CP$ score to identify the strongest correlates of persuasion (§6.3). To further understand the influence of the various potential predictors of an instance’s persuasion, we implement a logistic regression model on stubborn and persuaded instances. We take as dependent variables the popularity of the subject and object, the number of edits to the fact, and the semantic entropy before and after context is provided (§6.4). We standardise each parameter with a $z$-score transformation for interpretable comparison.
@@ 6.1 General Performance
Table 1 shows the percentage of stubborn and persuaded instances and Table 2 shows the performance of our three investigated models across the three partitions of our dataset with and without additional context. Semantic Entropy ($SE$) reflects the semantic variety of model output, which can indicate internal memory conflict (§4.2). Lower $SE$ can be interpreted as more consistent, and less conflicted, output. On the other hand, our Coherent Persuasion ($CP$) score quantifies the actual efficacy of the context in shifting an LM’s output distribution (§4.3).
@@
Accuracy is typically greatest with Llama-2 across all partitions. We also see that Llama-2 has the greatest degree of persuasion (in both persuasion score and % persuaded) across all partitions. However, Llama-2 also shows the greatest entropy across all instances. This suggests Llama-2 is typically the most effective at utilising context, yet also shows the greatest intra-memory conflict. Typically, entropy decreases with context; however, Mistral and Qwen2 show a slight increase in entropy for the Static partition of the dataset; this likely owes to the introduction of intra-memory conflict for instances with typically no competing objects. Furthermore, we typically see the greatest accuracy and persuasion (bolded) on the Static partition. In contrast, we typically see the lowest entropy (bolded) on the Temporal dataset. This is typically when context is provided, suggesting that the introduction of context reduces the exhibited intra-memory conflict for temporal facts.
@@
Our results are somewhat similar to contemporary work Fierro et al. (2024), who also find the greatest accuracy with Llama-2; however, they also find the highest confidence with this model, though they presumably use the output of softmax layer which is more vulnerable to syntactic uncertainties Kuhn et al. (2023). Furthermore, they find reduced update accuracy for Llama-2 relative to other models across mutable and immutable facts. In contrast, we surprisingly find that manipulations to static, immutable facts are most easily accepted by the model in context. In contrast, the temporal and disputable facts, which are expected to have greater variability in the model’s training dataset, and thus higher intra-memory conflict, have a greater proportion of stubborn instances across all models. We expect this stems from our differences in dataset collection, as Fierro et al. (2024) determine mutability via relationship type, whereas we verify the ‘mutability’ of our facts with our proxy measures. Ultimately, our results raise concerns for the efficacy of context usage in retrieval-augmented text generation, as we find that the most commonly adapted facts (dynamic facts) are the most difficult to update with context.
@@
We now investigate potential factors contributing to context usage (‘persuasion’) and neglect (‘stubbornness’). For the main paper, we present results for Llama-2, due to its high performance, on the temporal partition of the dataset, as this partition allows for easy comparison of the effect of dynamicity on model performance. We provide all the additional graphs for the remaining models and partitions in Appendix B.
Figure 2: The distribution of loss ($\mathcal{L}(a_{c},x_{i;q})$) for generating a particular answer ($a_{c}$) given only the question ($x_{i;q}$) for each partition of the dataset.
Figure 3: The instance-level relationship between Coherent Persuasion score (§4.3) and 3 possible factors that impact persuasion: semantic entropy, subject popularity and object temporality, alongside their Pearson correlation scores. Temporality shows the strongest relationship with persuasion. We also highlight two behaviours of interest: persuaded and stubborn instances.
6.2 Obstacles to Persuasion
To further investigate why dynamic facts show a greater proportion of stubborn instances yet a smaller divergence in output distributions, we look at the loss when we force the target answer ($a_{c}$) for the contextless input $x_{i;q}$. The loss reflects the likelihood of an output given the model’s trained parameters. A higher loss, thereby, also indicates greater change required to steer the LM to output the target answer. Thus, we expect higher loss on dynamic facts. Figure 2 shows the distribution of loss for each partition of the dataset. We construct the input only with the question and calculate the losses for all the tokens in the target answer ($a_{c}$). The losses from the tokens in the target answer are averaged for each question. In Figure 2, we observe that losses for temporal and disputable questions are greater than the loss for static questions; the significance is confirmed with a Welch’s $t$-test ($p=9.2e-64$, $p=3.4e-18$). Therefore, while static questions show a greater $CP$ score in Table 2, we find there is more effort involved in updating the LM’s parameters for dynamic facts, which may be unachievable with only additional context.
6.3 Interactors with Persuasion
@@
While popularity has been previously reported to negatively impact persuasion Du et al. (2024); Mallen et al. (2023), popular instances are not overly represented among the stubborn instances, and the strength of the correlation ($r(2494)=0.101,p=4e-7$), though statistically significant, is still weak. In contrast, temporality scores show the strongest correlation with $CP$ ($r(2944)=-0.27,p=1.6e-42$), suggesting there is a stronger relation between persuasion and object temporality than between persuasion and subject popularity or semantic entropy.
Furthermore, given the strong differences in the distributions in Figure 3, we can see that semantic entropy does not indicate epistemic uncertainty (which is approximated by subject popularity) nor aleatoric uncertainty (approximated by temporality) (though semantic entropy was previously used as a measure of aleatoric uncertainty in other work Gao et al. (2024)). This suggests that other factors may contribute to semantic entropy, outside of exclusively intra-memory conflict. While knowledge conflicts are a straightforward way to elicit meaningful semantic inconsistencies, they can come from other effects of model training (i.e., overlaps of objects in model memory space, spurious signals, common word co-occurrences), which may be picked up by the semantic entropy measure in addition to intra-memory conflict.